OpenStudy (anonymous):

What is the solution set of |x – 6| = 8?

OpenStudy (unklerhaukus):

There are two solutions

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Since absolute value takes negative things and makes them positive, this will be true if x-6 = 8 or if x-6 =-8

OpenStudy (unklerhaukus):

\[\qquad\qquad|x-6|=8\]\[\qquad\qquad\qquad\downarrow \]\[(x-6)=8\qquad\text{or} \qquad -(x-6)=8\]\[x-6+6=8+6\qquad\text{or} \qquad -x+6-6=8-6\]\[x=14\qquad\text{or} \qquad -x=2\]\[\qquad \qquad\text{or} \qquad x=-2\] \[x=\{-2,14\}\]

OpenStudy (anonymous):

thanks